OpenStudy (anonymous):

How do you graph 3x+5y=10?

OpenStudy (anonymous):

express y in terms of x

OpenStudy (anonymous):

Write this equation in y = mx + b form: 3x + 5y = 10 5y = 3x - 10 y = (3/5)x - 2 Now, the slope of this equation is (3/4) and the y intercept is b = -2. Slope = rise/run Therefore, plot the y intercept at (0, -2). And now move three units up (rise) and then 5 units to the right (run). And plot that point. Then continue on the same way until you have sufficient points to create a straight line.

OpenStudy (anonymous):

@ katragaddasaichandra .. Thank you SO MUCH. You explained it so well!! :)